MR has the following bank balance sheet
Assets
. Required reserves. X
.Excess reserves. R50 million


Liabilities
Deposits Y


The required reserves ratio is 10% .what is the value of MR bank's deposits from the given options.

I.R50 million
ii.R55.6 million
iii.R10 million
iv.R45 million
1
Expert's answer
2021-09-27T09:04:43-0400

We have:

Reserve ratio = 10% = 0.1

Excess reserve ratio = 1 - 0.1 = 0.9

"Deposits=\frac{ExcessReserves}{0.9}=\frac{50}{0.9}\\"

= 55.6

To equalize balance sheet, we have:

Answer (ii) R55.6 million